a rectangle is 3cm longer than it is wide. IF the length and width of the rectangle are each increased by 2cm, the area is increased by 54cm squared. What are the dimensions of the original rectangle?
Answer by mukhopadhyay(490)

Let the width be x cm
The length = (x+3)
Area=x(x+3)
Now, if the length and width are increased by 2cm, length=(x+2) and width=(x+5)
Area=(x+2)(x+5)
(x+2)(x+5)=x(x+3)+54
=>x^2+7x+10=x^2+3x+54
=>7x+10=3x+54
=>4x=44
=>x=11
Dimensions of the original rectangle is 14cm X 11 cm